In a normal population where the mean is 900 and the standard deviation is 25; what value of X is the 67th percentile?  
Answer #1

solution

Using standard normal table,

P(Z < z) = 67%

= P(Z < z) = 0.67

= P(Z < 0.44) = 0.67

z =0.44 ( Using standard normal z table,)

Using z-score formula  

x= z * +

x= 0.44*25+900

x=911